### v2

@@ 2.2 NLoS Channel
$\begin{split}g_{\scriptscriptstyle{k,n,\text{LED}}}^{\scriptscriptstyle\text{NLOS}}=\begin{cases}\frac{\rho\,(m+1)\,A}{2\pi\,(d_{\scriptscriptstyle{n,\text{LED}}}+d_{\scriptscriptstyle{k,n}})^{2}}\,\cos^{m}\!\bigl(\phi_{\scriptscriptstyle{n,\text{LED}}}\bigr)\,\cos\!\bigl(\psi_{\scriptscriptstyle{k,n}}\bigr)\,\mathbb{G}_{\scriptscriptstyle f}\,\mathbb{G}_{\scriptscriptstyle c}\,,\\ \hskip 80.00012pt\text{for }0\leq\psi_{\scriptscriptstyle{k,n}}\leq\Psi_{\scriptscriptstyle\text{FoV}}\,,\\ 0\,,\hskip 105.00015pt\text{otherwise}\end{cases}\end{split}$ (4)
where $d_{\scriptscriptstyle{n,\text{LED}}}$ and $d_{\scriptscriptstyle{k,n}}$ denote the Euclidean distances from the LED to element $n$ and from element $n$ to user $k$, respectively. The angles $\phi_{\scriptscriptstyle{n,\text{LED}}}$ and $\psi_{\scriptscriptstyle{k,n}}$ represent the irradiance and incidence angles associated with the path via element $n$, and the gain is limited by the receiver field of view $\Psi_{\scriptscriptstyle\text{FoV}}$ through the same filter and concentrator terms used in the LoS case.
This work considers realistic assumptions and accounts for the time delay in light propagation for both LoS and NLoS links. For user $k$, the propagation delay of the direct LoS path is $\tau_{\scriptscriptstyle{k,\text{LED}}}=d_{\scriptscriptstyle{k,\text{LED}}}/c$, while the delay associated with the path via element $n$ is $\tau_{\scriptscriptstyle{k,n,\text{LED}}}=(d_{\scriptscriptstyle n,\text{LED}}+d_{\scriptscriptstyle k,n})/c$, where $c$ is the speed of light. To represent the overall channel observed by user $k$, the LoS and NLoS contributions are combined into the CIR $q_{k}(t)$ as [chen2023freqdomainIRS]
@@
The signal-to-noise ratio (SNR) at frequency $f$ for user $k$ compares the power of the modulated signal to the noise power at the PD and is written as [chen2023freqdomainIRS]
@@ 3 Secrecy Performance Analysis
The starting point of the analysis is the achievable rate of a generic user $k$ in the DCO-OFDM link. Under the information theoretic model for a frequency selective channel, the rate $R_{k}$ is obtained by integrating over frequency [chen2023freqdomainIRS]
@@ 3.1 Secrecy Capacity for Non-Colluding Eavesdroppers
$C_{s}^{\text{non-coll}}=\max\left(0,R_{B}-\max_{j\in\{1..K\}}\{R_{E,j}\}\right)$ (10)
@@
For notational convenience, the LoS gain is denoted by $G_{\text{los},k}\triangleq g_{\scriptscriptstyle{k,\text{LED}}}^{\scriptscriptstyle\text{LOS}}$, the $i$th NLoS gain by $G_{i,k}\triangleq g_{\scriptscriptstyle{k,i,\text{LED}}}^{\scriptscriptstyle\text{NLOS}}$, the LoS delay by $\tau_{\text{los},k}\triangleq\tau_{\scriptscriptstyle{k,\text{LED}}}$, and the $i$th NLoS delay by $\tau_{i,k}\triangleq\tau_{\scriptscriptstyle{k,i,\text{LED}}}$. The frequency independent SNR prefix is written as $\Lambda=\frac{2T_{\scriptscriptstyle s}P_{\scriptscriptstyle\text{opt}}^{2}R_{\scriptscriptstyle\text{pd}}^{2}}{\Gamma\,\mathsf{k}^{2}\,\mathcal{N}}$, where a flat power distribution $E(f)=1$ is assumed.
By substituting (8) into (7) and then into (9), the rate integral for user $k$ can be written as
@@
To simplify this expression, we separate the frequency–independent term from the frequency–dependent term. The frequency–independent part is collected in
$D_{k}=1+\Lambda\left(G_{\text{los},k}^{2}+\sum_{i=1}^{N_{\text{irs}}}G_{i,k}^{2}\right)$ (12)
@@
$R_{k}=\underbrace{\frac{1}{2T_{\scriptscriptstyle s}}\log_{2}\left(D_{k}\right)}_{R_{k,1}}+\underbrace{\int_{0}^{\frac{1}{2T_{\scriptscriptstyle s}}}\log_{2}\left(1+\frac{\Lambda N_{k}(f)}{D_{k}}\right)df}_{R_{k,2}}$ (14)
@@
The integral term $R_{k,2}$ involves the logarithm of a sum of cosine functions and is not available in closed form. To obtain an analytical approximation, the following steps are used. First, the cosine terms in $N_{k}(f)$ are expressed in terms of the delay differences and a balanced approximation for $\cos(x)$ based on $\frac{\sin(x)}{x}$ is employed, inspired by the inequality in [Abramowitz1964], which establishes that $\cos(x)\leq\frac{\sin(x)}{x}\leq 1$ for $0\leq x\leq\pi$. Second, the main contribution of the integral is assumed to come from the dominant terms over the bandwidth $0,f_{\text{max}}]$ with $f_{\text{max}}=\frac{1}{2T_{s}}$. This procedure leads to the following approximation for the total rate $R_{k}$
@@
$\left.+\frac{1}{T_{\scriptscriptstyle s}}\sum_{i=1}^{N_{\text{irs}}}\sum_{u<i}G_{i,k}G_{u,k}\left(1-\frac{\sin\left(\frac{\pi(\tau_{u,k}-\tau_{i,k})}{T_{\scriptscriptstyle s}}\right)}{\frac{\pi(\tau_{u,k}-\tau_{i,k})}{T_{\scriptscriptstyle s}}}\right)\right)$ (15)
@@
$C_{s}^{\text{non-coll}}\approx\max\left(0,R_{B}^{\text{approx}}-\max_{j\in\{1..K\}}\{R_{E,j}^{\text{approx}}\}\right)$ (16)
@@ 3.2 Secrecy Capacity for Colluding Eavesdroppers
$C_{s}^{\text{coll}}=\left[\int_{0}^{\frac{1}{2T_{\scriptscriptstyle s}}}\log_{2}\left(\frac{1+\Lambda\,|Q_{B}(f)|^{2}}{1+\Lambda\sum_{j=1}^{K}|Q_{E_{j}}(f)|^{2}}\right)df\right]^{+}$ (23)
@@ 4.1 Secrecy Capacity Maximisation Problem
$Q_{k}(f,\mathbf{a})=g_{\scriptscriptstyle{k,\text{LED}}}^{\scriptscriptstyle\text{LOS}}e^{-j2\pi f\tau_{\scriptscriptstyle{k,\text{LED}}}}+\sum_{n\in\mathcal{N}_{k}(\mathbf{a})}g_{\scriptscriptstyle{k,n,\text{LED}}}^{\scriptscriptstyle\text{NLOS}}e^{-j2\pi f\tau_{\scriptscriptstyle{k,n,\text{LED}}}}$ (25)
@@ 4.2 Proposed PPO Based Solution
  • Agent: The RL agent is the IRS controller, which must learn an allocation policy for the reflecting elements.
  • State ($s_{t}$): The process is modelled as a sequential allocation over $N_{\text{irs}}$ decision steps. At decision step $t\in\{1,\dots,N_{\text{irs}}\}$, the agent allocates the $t$th element. The state $s_{t}$ contains the allocation history up to that point and is written as $s_{t}=(a_{1},a_{2},\dots,a_{t-1})$. The initial state is $s_{1}=\emptyset$. This representation allows the policy to capture dependencies between element allocations, where the decision for element $t$ can depend on earlier decisions.
  • Action ($a_{t}$): At each decision step $t$, the agent selects an action $a_{t}$ from the discrete action set $\mathcal{A}_{t}=\{B,E_{1},\dots,E_{K}\}$. This action assigns the $t$th IRS element to one of the users and therefore determines to which user that element contributes coherent NLoS power.
  • Policy ($\pi_{\theta}(a_{t}|s_{t})$): The agent’s policy is represented by a deep neural network with parameters $\theta$. The network takes the current state $s_{t}$ as input and outputs a probability distribution over all actions in $\mathcal{A}_{t}$. The action $a_{t}$ is sampled from this distribution. Training aims to adjust the parameters $\theta$ so that the resulting allocation sequences yield higher secrecy capacity.
  • Reward ($R$): The task is episodic. One episode consists of $N_{\text{irs}}$ decisions, which produce a complete allocation vector $\mathbf{a}=(a_{1},\dots,a_{N_{\text{irs}}})$. After the last element has been allocated, the environment returns a terminal reward
@@
The PPO algorithm is used to optimise the policy $\pi_{\theta}$ by iteratively collecting experience and updating the network parameters. To improve training stability and efficiency, we collect a batch of multiple episodes before performing policy updates. Each trajectory in the batch is of the form
$\tau=(s_{1},a_{1},\dots,s_{N_{\text{irs}}},a_{N_{\text{irs}}},R)$
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the probability ratio and $\hat{A}_{t}$ is an estimate of the advantage at step $t$, computed using the Generalized Advantage Estimation (GAE) method with parameters $\gamma=0.99$ and $\lambda=0.95$. The clipping operation limits the change in $r_{t}(\theta)$ and stabilises the policy updates. The value network parameters $\phi$ are updated by minimising the squared error loss $L^{V}=(V_{\phi}(s_{t})-R)^{2}$.
The overall training procedure is summarised in Algorithm 1. The algorithm records the best allocation vector $\mathbf{a}_{\text{best}}$ observed during training, which is used in the numerical evaluation.
@@
1:Input: Number of training episodes $N_{\text{ep}}$, batch size $N_{\text{batch}}$, clipping factor $\epsilon$, learning rates $\alpha_{\pi}$ and $\alpha_{V}$, discount factor $\gamma$, GAE parameter $\lambda$, channel gains $g^{\scriptscriptstyle\text{LOS}}_{\scriptscriptstyle k,\text{LED}}$ and $g^{\scriptscriptstyle\text{NLOS}}_{\scriptscriptstyle k,n,\text{LED}}$ for all $k$ and $n$, and fixed constants $\Lambda$, $T_{s}$, $K$, $N_{\text{irs}}$.
2:Output: Allocation vector $\mathbf{a}_{\text{best}}$ maximising (27).
@@
18:   Compute advantage estimates $\hat{A}_{t}$ for all steps in $\mathcal{D}$ using GAE($\gamma$, $\lambda$).
@@
Table 1: PPO Hyperparameters
@@ 5 Simulation Results
Fig. 2 shows the achievable rates of Bob, Eve1, and Eve2 as functions of the transmit optical power in the colluding eavesdropper scenario, for the best–case geometry where Bob is located directly beneath the LED and thus receives the strongest line–of–sight signal. The solid curves correspond to the LoS–only baseline, whereas the dashed curves represent the LoS+IRS case with PPO–based allocation. It can be observed that Bob’s rate is consistently higher when the IRS is used. At $P_{\text{opt}}=6$ W, the PPO–optimised allocation provides an increase of roughly 67% in Bob’s rate compared with the LoS–only link. In contrast, the rates of Eve1 and Eve2 are slightly reduced by the optimised allocation: at $P_{\text{opt}}=6$ W the rate of Eve1 decreases by about 5.8%, while that of Eve2 decreases by about 6.7% relative to their LoS–only values.

### v3

@@ 2.2 NLoS Channel
$\begin{split}g_{\scriptscriptstyle{k,n,\text{LED}}}^{\scriptscriptstyle\text{NLOS}}=\begin{cases}\frac{\rho\,(m+1)\,A}{2\pi\,(d_{\scriptscriptstyle{n,\text{LED}}}+d_{\scriptscriptstyle{k,n}})^{2}}\,\cos^{m}\!\bigl(\phi_{\scriptscriptstyle{n,\text{LED}}}\bigr)\,\cos\!\bigl(\psi_{\scriptscriptstyle{k,n}}\bigr)\,\mathbb{G}_{\scriptscriptstyle f}\,\mathbb{G}_{\scriptscriptstyle c}\,,\\ \hskip 80.00012pt\text{for }0\leq\psi_{\scriptscriptstyle{k,n}}\leq\Psi_{\scriptscriptstyle\text{FoV}}\,,\\ 0\,,\hskip 105.00015pt\text{otherwise},\end{cases}\end{split}$ (4)
where $d_{\scriptscriptstyle{n,\text{LED}}}$ and $d_{\scriptscriptstyle{k,n}}$ denote the Euclidean distances from the LED to element $n$ and from element $n$ to user $k$, respectively. The angles $\phi_{\scriptscriptstyle{n,\text{LED}}}$ and $\psi_{\scriptscriptstyle{k,n}}$ represent the irradiance and incidence angles associated with the path via element $n$, and the gain is limited by the receiver FoV, $\Psi_{\scriptscriptstyle\text{FoV}}$, through the same filter and concentrator terms used in the LoS case.
This work considers realistic assumptions and accounts for the time delay in light propagation for both LoS and NLoS links. For user $k$, the propagation delay of the direct LoS path is $\tau_{\scriptscriptstyle{k,\text{LED}}}=d_{\scriptscriptstyle{k,\text{LED}}}/c$, while the delay associated with the path via element $n$ is $\tau_{\scriptscriptstyle{k,n,\text{LED}}}=(d_{\scriptscriptstyle n,\text{LED}}+d_{\scriptscriptstyle k,n})/c$, where $c$ is the speed of light. To represent the overall channel observed by user $k$, the LoS and NLoS contributions are combined into the CIR $q_{k}(t)$ as [11]
@@
The signal-to-noise ratio (SNR) at frequency $f$ for user $k$ compares the power of the modulated signal to the noise power at the PD and is written as [11]
@@ 3 Secrecy Performance Analysis
The starting point of the analysis is the achievable rate of a generic user $k$ in the DCO-OFDM link. Under the information theoretic model for a frequency selective channel, the rate $R_{k}$ is obtained by integrating over frequency [11]
@@ 3.1 Secrecy Capacity for Non-Colluding Eavesdroppers
$C_{s}^{\text{non-coll}}=\max\left(0,R_{B}-\max_{j\in\{1..K\}}\{R_{E,j}\}\right),$ (10)
@@
For notational convenience, the LoS gain is denoted by $G_{\text{los},k}\triangleq g_{\scriptscriptstyle{k,\text{LED}}}^{\scriptscriptstyle\text{LOS}}$, the $i$th NLoS gain by $G_{i,k}\triangleq g_{\scriptscriptstyle{k,i,\text{LED}}}^{\scriptscriptstyle\text{NLOS}}$, the LoS delay by $\tau_{\text{los},k}\triangleq\tau_{\scriptscriptstyle{k,\text{LED}}}$, and the $i$th NLoS delay by $\tau_{i,k}\triangleq\tau_{\scriptscriptstyle{k,i,\text{LED}}}$. The frequency independent SNR prefix is written as $\Lambda=\frac{2T_{\scriptscriptstyle s}P_{\scriptscriptstyle\text{opt}}^{2}R_{\scriptscriptstyle\text{pd}}^{2}}{\Gamma\,\mathsf{k}^{2}\,\mathcal{N}}$, where a flat power distribution $E(f)=1$ is assumed. By substituting (8) into (7) and then into (9), the rate integral for user $k$ can be written as
@@
To simplify this expression, we separate the frequency-independent term from the frequency–dependent term. The frequency–independent part is collected in
$D_{k}=1+\Lambda\left(G_{\text{los},k}^{2}+\sum_{i=1}^{N_{\text{irs}}}G_{i,k}^{2}\right),$ (12)
@@
$R_{k}=\underbrace{\frac{1}{2T_{\scriptscriptstyle s}}\log_{2}\left(D_{k}\right)}_{R_{k,1}}+\underbrace{\int_{0}^{\frac{1}{2T_{\scriptscriptstyle s}}}\log_{2}\left(1+\frac{\Lambda N_{k}(f)}{D_{k}}\right)df}_{R_{k,2}},$ (14)
@@
The integral term $R_{k,2}$ involves the logarithm of a sum of cosine functions and is not available in closed form. To obtain an analytical approximation, the following steps are used. First, the cosine terms in $N_{k}(f)$ are expressed in terms of the delay differences and a balanced approximation for $\cos(x)$ based on $\frac{\sin(x)}{x}$ is employed, inspired by the inequality in [2], which establishes that $\cos(x)\leq\frac{\sin(x)}{x}\leq 1$ for $0\leq x\leq\pi$. Second, the main contribution of the integral is assumed to come from the dominant terms over the bandwidth $0,f_{\text{max}}]$ with $f_{\text{max}}=\frac{1}{2T_{s}}$. This procedure leads to the following approximation for the total rate $R_{k}$
@@
$\left.+\frac{1}{T_{\scriptscriptstyle s}}\sum_{i=1}^{N_{\text{irs}}}\sum_{u<i}G_{i,k}G_{u,k}\left(1-\frac{\sin\left(\frac{\pi(\tau_{u,k}-\tau_{i,k})}{T_{\scriptscriptstyle s}}\right)}{\frac{\pi(\tau_{u,k}-\tau_{i,k})}{T_{\scriptscriptstyle s}}}\right)\right).$ (15)
@@
$C_{s}^{\text{non-coll}}\approx\max\left(0,R_{B}^{\text{approx}}-\max_{j\in\{1..K\}}\{R_{E,j}^{\text{approx}}\}\right),$ (16)
@@ 3.2 Secrecy Capacity for Colluding Eavesdroppers
$C_{s}^{\text{coll}}=\left[\int_{0}^{\frac{1}{2T_{\scriptscriptstyle s}}}\log_{2}\left(\frac{1+\Lambda\,|Q_{B}(f)|^{2}}{1+\Lambda\sum_{j=1}^{K}|Q_{E_{j}}(f)|^{2}}\right)df\right]^{+},$ (23)
@@ 4.1 Secrecy Capacity Maximisation Problem
$Q_{k}(f,\mathbf{a})=g_{\scriptscriptstyle{k,\text{LED}}}^{\scriptscriptstyle\text{LOS}}e^{-j2\pi f\tau_{\scriptscriptstyle{k,\text{LED}}}}+\sum_{n\in\mathcal{N}_{k}(\mathbf{a})}g_{\scriptscriptstyle{k,n,\text{LED}}}^{\scriptscriptstyle\text{NLOS}}e^{-j2\pi f\tau_{\scriptscriptstyle{k,n,\text{LED}}}},$ (25)
@@ 4.2 Proposed PPO Based Solution
  • Agent. The RL agent is the IRS controller, which must learn an allocation policy for the reflecting elements.
  • State $s_{t}$. The process is modelled as a sequential allocation over $N_{\text{irs}}$ decision steps. At decision step $t\in\{1,\dots,N_{\text{irs}}\}$, the agent allocates the $t$th element. The state $s_{t}$ contains the allocation history up to that point and is written as $s_{t}=(a_{1},a_{2},\dots,a_{t-1})$. The initial state is $s_{1}=\emptyset$. This representation allows the policy to capture dependencies between element allocations, where the decision for element $t$ can depend on earlier decisions.
  • Action $a_{t}$. At each decision step $t$, the agent selects an action $a_{t}$ from the discrete action set $\mathcal{A}_{t}=\{B,E_{1},\dots,E_{K}\}$. This action assigns the $t$th IRS element to one of the users and therefore determines to which user that element contributes coherent NLoS power.
  • Policy $\pi_{\theta}(a_{t}|s_{t})$. The agent’s policy is represented by a deep neural network with parameters $\theta$. The policy network takes the state vector $s_{t}$ of dimension $N_{\text{irs}}+9$ as input. This input consists of the allocation history combined with user coordinates. The network processes this input through a hidden layer with $d_{h}=256$ neurons and ReLU activation. The output layer uses softmax activation to produce a probability distribution over $K+1$ possible actions. The action $a_{t}$ is sampled from this distribution. Training aims to adjust the parameters $\theta$ so that the resulting allocation sequences yield higher secrecy capacity.
  • Reward $R$. The task is episodic. One episode consists of $N_{\text{irs}}$ decisions, which produce a complete allocation vector $\mathbf{a}=(a_{1},\dots,a_{N_{\text{irs}}})$. After the last element has been allocated, the environment returns a terminal reward
@@
A separate value network with identical architecture is used to estimate the state value $V_{\phi}(s_{t})$ for advantage computation. This critic network also has input dimension $N_{\text{irs}}+9$ and a hidden layer with 256 neurons and ReLU activation. The output layer consists of a single linear neuron that produces the value estimate. The PPO algorithm is used to optimise the policy $\pi_{\theta}$ by iteratively collecting experience and updating the network parameters. To improve training stability and efficiency, we collect a batch of multiple episodes before performing policy updates. Specifically, experience is accumulated over $N_{\text{batch}}=64$ complete episodes before any network update is performed. Each trajectory in the batch is of the form
$\tau=(s_{1},a_{1},\dots,s_{N_{\text{irs}}},a_{N_{\text{irs}}},R),$
@@
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the probability ratio and $\hat{A}_{t}$ is an estimate of the advantage at step $t$, computed using the Generalized Advantage Estimation method with parameters $\gamma=0.99$ and $\lambda=0.95$. The clipping operation limits the change in $r_{t}(\theta)$ and stabilises the policy updates. The actor and critic networks are updated with learning rates $\alpha_{\pi}=1\times 10^{-4}$ and $\alpha_{V}=5\times 10^{-4}$ respectively. The value network parameters $\phi$ are updated by minimising the squared error loss $L^{V}=(V_{\phi}(s_{t})-R)^{2}$.
The complete set of hyperparameters is summarised in Table 1 to ensure reproducibility of the proposed approach. The training proceeds for $N_{\text{ep}}=1500$ episodes with the batch update schedule performed every 64 episodes as described above. The overall training procedure is summarised in Algorithm 1. The algorithm records the best allocation vector $\mathbf{a}_{\text{best}}$ observed during training, which is used in the numerical evaluation.
The computational complexity of the IRS allocation problem motivates the use of reinforcement learning rather than exhaustive search. For $K$ eavesdroppers and $N_{\text{irs}}$ elements, each element can be allocated to one of $K+1$ users, resulting in $(K+1)^{N_{\text{irs}}}$ possible allocations. Evaluating the secrecy capacity for each allocation requires numerical integration over the signal bandwidth, making exhaustive enumeration computationally prohibitive. As the number of eavesdroppers or IRS elements increases, the search space grows exponentially. The proposed PPO-based approach circumvents this exponential complexity by learning a policy through episodic training with $N_{\text{ep}}=1500$ episodes, where the training complexity scales linearly as $O(N_{\text{ep}}\cdot N_{\text{irs}}\cdot(K+1))$. Once trained, the policy network generates allocations through simple forward passes with $O(1)$ complexity per inference, enabling practical reconfiguration when channel conditions change due to user mobility. The linear scaling with respect to $N_{\text{irs}}$ and $K$ makes the approach feasible even for larger systems, unlike exhaustive search, which grows exponentially.
@@
1:Input. Hyper-parameters as per Table 1. Channel gains $g^{\scriptscriptstyle\text{LOS}}_{\scriptscriptstyle k,\text{LED}}$ and $g^{\scriptscriptstyle\text{NLOS}}_{\scriptscriptstyle k,n,\text{LED}}$ for all $k$ and $n$. Fixed constants $\Lambda$, $T_{s}$, $K$, $N_{\text{irs}}$.
2:Output. Allocation vector $\mathbf{a}_{\text{best}}$ maximising (27).
@@
18:   Compute advantage estimates $\hat{A}_{t}$ for all steps in $\mathcal{D}$ using GAE with $\gamma$ and $\lambda$.
@@
Table 1: PPO Hyper-parameters for Reproducible Training Configuration.
@@ 5 Simulation Results
Fig. 2 shows the achievable rates of Bob, Eve1, and Eve2 as functions of the transmit optical power in the colluding eavesdropper scenario, for the best–case geometry where Bob is located directly beneath the LED and thus receives the strongest LoS signal. The solid curves correspond to the LoS–only baseline, whereas the dashed curves represent the LoS+IRS case with PPO–based allocation. It can be observed that Bob’s rate is consistently higher when the IRS is used. At $P_{\text{opt}}=6$ W, the PPO–optimised allocation provides an increase of roughly 67% in Bob’s rate compared with the LoS–only link. In contrast, the rates of Eve1 and Eve2 are slightly reduced by the optimised allocation: at $P_{\text{opt}}=6$ W the rate of Eve1 decreases by about 5.8%, while that of Eve2 decreases by about 6.7% relative to their LoS–only values.
@@
Fig. 9 validates the closed-form approximate rate expression in (15) against the exact numerical integration in (9) for two different IRS configurations. The results confirm that the analytical approximation remains highly consistent with the numerical values across the entire range of transmit power for a representative edge user. As shown in the figure, increasing the number of IRS elements results in a higher achievable rate. This close agreement between the curves demonstrates that the proposed analytical model accurately captures the channel behaviour while providing a computationally efficient alternative to numerical integration.
Figure 9: Validation of the analytical rate expression against exact numerical integration for different numbers of IRS elements.